OpenStudy (anonymous):

Find the next term of the sequence: 512 256 –512 –256

OpenStudy (anonymous):

\[Sequence >> \frac{ 1 }{ 128 },-\frac{ 1 }{ 16 },\frac{ 1 }{ 2 }, -4,32,...\] @satellite73

OpenStudy (anonymous):

looks like you are multiplying by 8?

OpenStudy (anonymous):

oh scratch that, by \(-8\)

OpenStudy (anonymous):

if it is not clear, take any ratio\[\frac{32}{-4}=-8\]

OpenStudy (anonymous):

\[\frac{-4}{\frac{1}{2}}=-8\] etc

OpenStudy (anonymous):

So -256

sammixboo (sammixboo):

Yes, because 32 x -8 = -256 :)